Prev Next

SW&IT Meeting Minutes 2019-09-02

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20190826
LR20190826 - Bec, Matthieu 26.0
LR20190826 - Holmberg, Carl 0.0
LR20190826 - Colson, Andrew 32.0
LR20190826 - Edmeades, David 39.0
LR20190826 - Sola, Igor 0.0
LR20190826 - Kubanek, Petr 33.5
LR20190826 - Hooper, Stephen 72.0
LR20190826 - Zhang, Yang 40.0
LR20190826 - Zhang, Xianyu 48.0

News from management

Previous Action Items

  • explore pros/cons of Timestamps vs. uint64_t for timestamps in MongoDb (CH, YZ)
  • ALH and AO (MB, XZ)
  • Luci clients on new obs machine (IS, SH)
  • colocation update, test connection with a laptop or some other linux computer, not dms (DE)
  • irtc on obs1 is running ice-3.7, confirm we're ok (PK)
  • create a github issue in ao-supervisor about the MTU setting (XZ)
  • FACSUM jump fix before restart (AC)
  • Azcamserver file transfer woes (AC, DE, MB)
  • update obswsup document with Amali LBTI input (MB) DONE

Open topics

  • wiki died again, let's restart it daily.

Team reports

PK

  • OSS / M2 SX, AGW troubleshooting (M2, learned how to use colours).
  • discovered std::out_of_range, std::move & move constructors
  • FITSpp work (FITS headers, saving & reading images)
  • PMC monitoring
  • re-enabled PMC & MCS telemetry
  • +a prints char as number in std::stream
  • GitLab + gitlab-runner, .gitlab-ci.yml

MB

  • OC, WPC, Management
  • DMS mongodb adapter
  • DMS alta predictions
  • DMS m1 surface (WIP)
  • Met with Amali about LBTI
next
  • Restart
  • Alta meeting
  • DMS

DE

  • Display jump narrowed down to display, sent 2 sample models for replacement
  • MOU signed by Gene; hardware procurement in progress
  • VM created on mountain, modifying KS file
  • Repaired Macbook returned; turnaround time is faster than anticipated in case anyone else needs a repair
  • Tyler Autodesk software sorted; Vault Pro needs to be installed last
  • Robert sent his PDF from his home laptop; no virus trigger. I'm confident that it was a false positive
  • RAM upgrades for stock workstations
  • VMware rep on vacation until 3 Sep; will make sure that I get the support sorted this week
  • Juan Pablo accounts
  • 6-core wiki.lbto seems to be happier; unsure if that is causal

SH

  • open LUCI script (window placement, etc.)
  • ARGOS bad drive (open new drive)
  • sw0401, sw0206 stacks repaired
  • mountain NUCs upgraded
  • Chris Bolyard's laptop installed
  • Chromium for kiosk mode switches determined for newest NUC install
  • Power up on Wednesday
  • Intro to room 507 beta installations (updated kickstart yet again)
  • support call on Sunday night... weird PMCR stuff, DGH opened issue.
  • upgraded Zabbix "monitor.mountain.lbto.org" to latest.
next:
  • kick start remote observing VM
  • VNC view on new rm507 computers
  • LBTI scripts
  • browser extension for bookmarks.

YZ

  • LBC/MODS:
    • Discussed with Olga for MODS ISIS and LBC telemetry requirements.
    • Create Python program to injest MODS ISIS telemetry data.
    • Create Python program to injest LBC housekeeping data.
    • Investigate methods related to LBC skyflat integration.
    • Gather some performance data related to LBC FPIA/WRC code.
  • AO-Cravat:
    • Working on a draft POI document.
    • Investigate chart.js for the possibility of plots interaction, resizing, zooming, etc.
    • Need to schedule a discussion with Steve Ertel for his Strehl code (request sent on Thursday but haven't heard from him yet).
  • Completed some safty training courses.

CH

  • Last week:
    • Refactored iMonnit script as class, added Mongo date query to avoid data download overlap.
    • Installed web-dms and nginx infrastructure on workstation.
    • Began work on dms-monnit backend to create JSON blob, querying Mongo rather than HDF5 files.
  • This week:
    • Finish dms-monnit backend.
    • Add to Polymer-based viewer, if needed.

XZ

Last week:
  • Support the ASM restart
  • Put git on the SX AdSec and DX AdSec
  • Review a WFS science paper
  • Talk to AO operation people for AO checkput
  • AO support
  • Help LBTI WFS installation and configuration
  • DX AdSec bump issue
This week:
  • AO support
  • AO telemetry requirement document

IS

Last week
  • PIT
    • release 2020A
  • OT
    • Adding proposal ID (ongoing)
  • Luci
    • Finish scripting of LUCI client start apps. The same scripting should work in Linux 32/64 and Mac
    • Started working on a command line option for taking grating calibration for Dave T.

This week:
  • Hot Wiring the Transient Universe VI conference presentation
  • Luci
    • Finish command line option for taking grating calibration for Dave T.
  • OT
    • Providing a table with filter capabilities will all the proposals Ids, pi names and emails
    • Fixing some of the MODS scripting issues
    • other bugs
Topic revision: r16 - 06 Sep 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback